10. workers: Larry, Moe, and Curly. Each works 10 hours a day and can produce two services: mowing lawns and washing Using the blue points (circle symbol), graph the production possibiities frontler (PPF) for this economy on the folowing graph. Then use the black the purpie point (diamond symbol) to identify point D on the graph.
mows lawns. (D) Using the blue points (circle symbol), graph the production possiblities frontier (PPF) for this economy on the folowing graph. Then use the black point (olus symbol) to identily point A, the green point (triangle symbol) to ldlentily point B, the orange point (square symbol) to identily point C and the purple point (diamond symbol) to identify point D on the graph. 0 5 15 20 2303 40 45 5 Quantity of Cars Washed frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because True or False: The production possibi each worker faces a constant trade-off between mowing lawns and washing cars. O True False e whether each of the
15 10+ 0 5 10 15 20 25 30 35 40 45 50 Quantity of Cars Washed True or False: The production possibilities frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because each worker faces a constant trade-off between mowing lawns and washing cars. O True O False of the following allocations is efficient or inefficient.
Answer #1

each worker works for 10 hours a day

lawns (in 1 hour ) cars washed ( in 1 hour )
lady 1 1
Moe 1 2
Curly 2 1
scenario lawns mowed cars washed
all three spend all Time mowing lawn (A) 10 + 10 + 20 = 40 0
all three spend all Time washing cars (B) 0 10 + 20 + 10 = 40
all three spend half of their time n each activity 5+ 5+ 10= 20 5+ 10 + 5 = 20
lary spends half his time on each , moe only washes and curly only mows lawns 5+ 20= 25 5 + 20 = 25

line connecting ADB is the production .frontier

5 5 to 20 25 3o 35 40 45 5 au washed

false ,

the production frontier is not a straight line.It is a smoothly bowed outwards curve .

allocation efficient inefficient
A yes
B yes
C inefficient
D yes